OpenStudy (anonymous):

Find the domain and the range of the function h(x)=sqrt(9-x^2) answer using interval notation please

OpenStudy (anonymous):

for domain solve \[9-x^2\geq 0\] or note that this function represents the upper half of a circle with radius 3 centered at the origin

OpenStudy (anonymous):

if that is not clear, start with \[x^2+y^2=3^2\] the entire circle solve for \(y\) and get \[y=\pm\sqrt{9-x^2}\] since you don't have the \(\pm\) part this is the upper half of the circle

OpenStudy (anonymous):

that makes the domain \([-3,3]\)